# Scaling Relations of Spectrum Form Factor and Krylov Complexity at Finite Temperature

### v1

@@ IV Scaling relation of GOE
In order to eliminate the size-dependent effects, a scaled ergodicity indicator denoted as $g^{\rm(GOE)}_{\rm S}(T)$ has been proposed. It remains invariant with respect to different system size. Notably, for a particular system that can be simulated by GOE, the spectrum spans approximately from $-\mathcal{E}\sqrt{L}$ to $\mathcal{E}\sqrt{L}$. We postulate that temperature should be scaled by a factor of $\sqrt{L}$. Using the definition of Heisenberg time, we find that it is system size-dependent, i.e., $t_{\rm H}=2\pi L/(E_{\rm max}-E_{\rm min})\sim 2\pi\sqrt{L}/\mathcal{E}$. This contrasts with the Thouless time. As such, we can shift the value of the ergodicity indicator by amount of $\log_{10}(\sqrt{L})$, and define scaled ergodicity indicator as
@@
Then the relation between $g^{\rm(GOE)}_{\rm S}(T)$ and $\alpha(T)$ can be well defined. It has been demonstrated that the ratio of the time it takes for the SFF to reach the minimum and the plateau is proportional to $\sqrt{L}/(1+\beta^{2}/2)$ for GUE Cotler et al. (2017b). Noticing that the SFF at Thouless time is close to the minimum and the Heisenberg time is pinned down by the presence of plateau, we conjecture the ratio of Heisenberg time and Thouless time for GOE has similar form and can be written as
$\frac{t_{\rm H}}{t_{\rm Th}}\propto\left(\frac{\sqrt{L}}{1+\beta^{2}\mathcal{E}^{2}/2}\right)^{\delta}.$ (13)
@@
where we have used $\beta=1/k_{\rm B}T$ and $C$ is a fitting parameter. In Fig. 3(b), we show that Eq.(14) indeed holds. For different system sizes, the curves collapse to a single one. By fitting Eq.(14) to numerical results, we find $\delta=0.25$ and $C=-0.21$. The origin of the error bars is the same as that for scaled ergodicity indicators.
@@ V Scaling relation of random spin model
Up to now, we work on the random matrices which describe sufficient complicated and chaotic systems. In this section, we study a 1D Ising model in transverse and longitudinal field, TL-Ising model in short. In our calculation, the longitudinal field takes a random value on each site. This model is nonintegrable and chaotic at uniform transverse and random longitudinal field Noh (2021); Nivedita et al. (2020); Trigueros and Lin (2022). And the energy spectrum follows the Wigner-Dyson distribution, which is a hallmark of quantum chaos. Besides, we choose the random parameter far away from the many-body-localization phase Trigueros and Lin (2022); De Tomasi et al. (2021) to reserve the chaotic characteristics. In this system, the linear scaling relation is tested.
@@
where $J$ is the near-neighbor coupling and taken as the energy dimension in the following discussion, $h$ represents the uniform transverse field and $g_{i}$ represents the site-dependent longitudinal field, $N$ is the total number of spin and $\hat{\sigma}_{i}^{a}$ with $a=x,z$ is the Pauli operator on $i$-th site. Formally, this operator should be considered in version of direct product that $\hat{\sigma}_{i}^{a}=...\otimes\underbrace{\hat{\mathbb{I}}_{2}}_{i-1\text{-thsite}}\otimes\underbrace{\hat{\sigma}^{a}}_{i\text{-th site}}\otimes\underbrace{\hat{\mathbb{I}}_{2}}_{i+1\text{-th site}}\otimes...$ with $\hat{\mathbb{I}}_{2}$ being the $2\times 2$ identity operator. When the longitudinal field is turned off ($g_{i}=0$ for each site), this model can be mapped to a free fermion model and becomes integrable. As shown in Noh (2021) with sightly applying uniform longitudinal field, the integrability is broken and the system tends to be chaotic. In our case, we take site-dependent longitudinal field $g_{i}$, which is sampled from a uniform distribution.
@@
The SFF of TL-Ising model at finite temperature is shown in Fig. 4(a), where $N=8$, $h=J$ and $g_{i}$ is sampled from $[-J,J]$. Each colored curve of SFF data is taken average from 200 samples. The “dip-ramp-plateau” structure can be observed, which confirms the existence of chaos. Using the same method illustrated in GOE, the Thouless time $t_{\rm Th}$ and Heisenberg time $t_{\rm H}$ are extracted from the SFF data. The Heisenberg time still serves as a time unit. At lower temperature, the SFF of TL-Ising model takes longer time to adhere to the analytical SFF of GOE, indicating larger Thouless time. Then the ergodicity indicator defined by Eq.(10) can be calculated. As expected, the ergodicity indicator shows a size dependence, seen in the inset of Fig. 5(a). As such, we need to scale the SFF to avoid the size effect.
The scaling method on TL-Ising model is different from the case of random matrix. For energy of TL-Ising model ranges from $\sim-JN$ to $\sim JN$. The temperature $k_{B}T/J$ is scaled to $k_{B}T/NJ$. In contrast to GOE, the Thouless time is proportion to $N^{2}$ Sierant et al. (2020); Šuntajs et al. (2020); D’Alessio et al. (2016). Therefore, the scaled ergodicity indicator for TL-Ising model is written as
@@
On the Krylov complexity side, we also use the Wightman inner product in the calculation of Lanczos coefficient $b_{n}$. In our calculation, the operator is chosen as $\hat{\sigma}^{(1)}_{z}\otimes\Pi_{i=2}^{N}\hat{\mathbb{I}}_{2}^{(i)}$. As the system is one dimensional, the linear growth of $b_{n}$ is corrected by logarithmic correction $b_{n}=A(T)n/{\rm ln}(n)$. Here $A(T)$ is the slope of $\left\{b_{n}\right\}$ sequence after logarithmic correction. Considering the disordered longitudinal field, we obtain the final $b_{n}$ by the mean value of 10 samples, and fit the results by a linear function. The temperature dependence of $A(T)$ is shown in Fig. 4(b). The points in figure represent the average of $A(T)$ from five repetitions and the error bars represent their standard deviations. It is obvious that $A(T)$ tends to saturate at high temperature limit, which indicates the TL-Ising model is chaotic but not sufficiently chaotic.
Then the scaling relation between the scaled ergodicity indicator and the Lanczos coefficients slope can be established. Inspired by the similarity of spectral properties between GOE and TL-Ising model Sierant et al. (2020); Šuntajs et al. (2020), we expect that the scaling law established in GOE should also hold for the TL-Ising model. Therefore, the counterpart of Eq. (14) can be written as
@@
As shown in Fig.5(b), for different system size, the linear scaling relation appears at relatively low temperature, which is consistent with the result of GOE. By fitting the numerical results to Eq.(17), we find $\delta=1.07$ and $C=2.94$ for TL-Ising model. We suspect that the deviation from linearity can be attributed to the property that this model is not sufficiently chaotic.
@@
In summary, we numerically investigate the finite-temperature behavior of K-complexity and SFF in the context of GOE and TL-Ising model. Within the GOE, the Lanczos sequence $\left\{b_{n}\right\}$ defined by the Wightman inner product exhibits notable linear growth at finite temperatures, which is not observed in the $\left\{b_{n}\right\}$ defined by the standard inner product. Utilizing the Wightman inner product, the slope $\alpha$ of $b_{n}$ saturates the universal upper bound $\pi k_{B}T$, indicating that the random matrix effectively simulates highly complex or chaotic systems. In the case of TL-Ising model, the behavior of $b_{n}$ defined by the Wightman inner product displays linear growth after a logarithmic correction. The slope $A$ converges to a specific value at infinite temperature, suggesting that the chaotic nature of this spin system is comparatively subdued compared to the GOE.

### v2

@@ IV Scaling relation of GOE
In order to eliminate the size-dependent effects, a scaled ergodicity indicator denoted as $g^{\rm(GOE)}_{\rm S}(T)$ has been proposed. It remains invariant with respect to different system size. Notably, for a particular system that can be simulated by GOE, the spectrum spans approximately from $-\mathcal{E}\sqrt{L}$ to $\mathcal{E}\sqrt{L}$. We postulate that temperature should be scaled by a factor of $\sqrt{L}$. Using the definition of Heisenberg time, we find that it is system size dependent, i.e., $t_{\rm H}=2\pi L/(E_{\rm max}-E_{\rm min})\sim 2\pi\sqrt{L}/\mathcal{E}$. This contrasts with the Thouless time. As such, we can shift the value of the ergodicity indicator by amount of $\log_{10}(\sqrt{L})$, and define scaled ergodicity indicator as
@@
Then the relation between $g^{\rm(GOE)}_{\rm S}(T)$ and $\alpha(T)$ can be well defined. It has been demonstrated that the ratio of the time it takes for the SFF to reach the minimum and the plateau is proportional to $\sqrt{L}/(1+\beta^{2}\mathcal{E}^{2}/2)$ for GUE Cotler et al. (2017b). Noticing that the SFF at Thouless time is close to the minimum and the Heisenberg time is pinned down by the presence of plateau, we conjecture the ratio of Heisenberg time and Thouless time for GOE has similar form and can be written as
$\frac{t_{\rm H}}{t_{\rm Th}}\propto\sqrt{L}\left(\frac{1}{1+\beta^{2}\mathcal{E}^{2}/2}\right)^{\delta}.$ (13)
@@
where we have used $\beta=1/k_{\rm B}T$ and $C$ is a fitting parameter. In Fig. 3(b), we show that Eq.(14) indeed holds around the temperature range of $k_{\rm B}T/{\cal E}\in[0.1,2]$. For different system sizes, the curves collapse to a single one. By fitting Eq.(14) to numerical results, we find $\delta=0.25$ and $C=-0.26$. The origin of the error bars is the same as that for scaled ergodicity indicators.
@@ V Scaling relation of random spin model
Up to now, we work on the random matrices ensemble which describes the sufficiently complicated and chaotic systems. In this section, we study a 1D Ising model in transverse and longitudinal field, TL-Ising model in short. In our calculation, the longitudinal field takes a random value on each site. This model is nonintegrable and chaotic at uniform transverse and random longitudinal field Noh (2021); Nivedita et al. (2020); Trigueros and Lin (2022). And the energy spectrum follows the Wigner-Dyson distribution, which is a hallmark of quantum chaos. Besides, we choose the random parameter far away from the many-body-localization phase Trigueros and Lin (2022); De Tomasi et al. (2021) to preserve the chaotic characteristics. In this system, the linear scaling relation is tested.
@@
where $J$ is the near-neighbor coupling and taken as the energy unite in the following discussion, $h$ represents the uniform transverse field and $g_{i}$ represents the site-dependent longitudinal field, $N$ is the total number of spin and $\hat{\sigma}_{i}^{a}$ with $a=x,z$ is the Pauli operator on $i$-th site. Formally, this operator should be considered in version of direct product that $\hat{\sigma}_{i}^{a}=...\otimes\underbrace{\hat{\mathbb{I}}_{2}}_{i-1\text{-thsite}}\otimes\underbrace{\hat{\sigma}^{a}}_{i\text{-th site}}\otimes\underbrace{\hat{\mathbb{I}}_{2}}_{i+1\text{-th site}}\otimes...$ with $\hat{\mathbb{I}}_{2}$ being the $2\times 2$ identity operator. When the longitudinal field is turned off ($g_{i}=0$ for each site), this model can be mapped to a free fermion model and becomes integrable. As shown in Noh (2021), the integrability is broken in the presence of uniform longitudinal field and the system tends to be chaotic. In our case, we take site-dependent longitudinal field $g_{i}$, which is sampled from a uniform distribution.
@@
The SFF of TL-Ising model at finite temperature is shown in Fig. 4(a), where $N=8$, $h=J$ and $g_{i}$ is sampled from $[-J,J]$. Each colored curve of SFF data is taken average from 200 samples. The “dip-ramp-plateau” structure is observed, which confirms the existence of chaos. Using the same method illustrated in GOE, the Thouless time $t_{\rm Th}$ and Heisenberg time $t_{\rm H}$ are extracted from the SFF data. The Heisenberg time still serves as a time unit. At lower temperature, the SFF of TL-Ising model takes longer time to adhere to the analytical SFF of GOE, indicating larger Thouless time. Then the ergodicity indicator defined by Eq.(10) can be calculated. As expected, the ergodicity indicator shows a size dependence, seen in the inset of Fig. 5(a). As such, we need to scale the SFF to avoid the size effect.
The scaling method on TL-Ising model is different from the case of random matrix. For energy of TL-Ising model ranges from $\sim-JN$ to $\sim JN$. The temperature $k_{B}T/J$ is scaled to $k_{B}T/NJ$. In contrast to GOE, the Thouless time is proportion to $N^{2}$ Sierant et al. (2020); Šuntajs et al. (2020). Therefore, the scaled ergodicity indicator for TL-Ising model is written as
@@
On the Krylov complexity side, we also use the Wightman inner product in the calculation of Lanczos coefficient $b_{n}$. In our calculation, the operator is chosen as $\hat{\sigma}^{(1)}_{z}\otimes\Pi_{i=2}^{N}\hat{\mathbb{I}}_{2}^{(i)}$. As the system is one dimensional, the linear growth of $b_{n}$ is corrected by logarithmic correction $b_{n}=A(T)n/{\rm ln}(n)$. Here $A(T)$ is the slope of $\left\{b_{n}\right\}$ sequence after logarithmic correction Parker et al. (2019). Considering the disordered longitudinal field, we obtain the final $b_{n}$ by the mean value of 10 samples, and fit the results by a linear function. The temperature dependence of $A(T)$ is shown in Fig. 4(b). The points in figure represent the average of $A(T)$ from five repetitions and the error bars represent their standard deviations. It is obvious that $A(T)$ tends to saturate at high temperature limit, which indicates the TL-Ising model is chaotic but not sufficiently chaotic.
Then the scaling relation between the scaled ergodicity indicator and the Lanczos coefficients slope is established. Inspired by the similarity of spectral properties between GOE and TL-Ising model Sierant et al. (2020); Šuntajs et al. (2020), we expect that the scaling law established in GOE should also hold for the TL-Ising model. Therefore, the counterpart of Eq. (14) can be written as
@@
As shown in Fig.5(b), for different system size, the linear scaling relation appears at relatively low temperature, which is consistent with the result of GOE. By fitting the numerical results to Eq.(17), we find $\delta=1.07$ and $C=2.94$ for TL-Ising model. As temperature $k_{\rm B}T/J>0.8$, a deviation from linearity arises. We suspect that the deviation is attributed to the property that this model is not sufficiently chaotic, leading to a saturation of both the spectral form factor (SFF) and the slope of Lanczos coefficients at finite temperatures, as illustrated in Fig.4. One of the side effects on SFF of keeping increasing temperature is the emergence of oscillations Cotler et al. (2017b), which causes the numerical SFF of Ising model to intersect with the analytical SFF of the GOE at an earlier time. Consequently, these oscillations further suppress the Thouless time and increase the value of scaled ergodicity indicator, which accounts for the observed deviation.
@@
In summary, we numerically investigate the finite-temperature behavior of K-complexity and SFF in the context of GOE and TL-Ising model. Within the GOE, the Lanczos coefficients $\left\{b_{n}\right\}$ defined by the Wightman inner product exhibit notable linear growth at finite temperatures, which is not observed in the $\left\{b_{n}\right\}$ defined by the standard inner product. Utilizing the Wightman inner product, the slope $\alpha$ of $b_{n}$ saturates the universal upper bound $\pi k_{B}T$, aligning with the fact that the random matrix effectively simulates highly complex or chaotic systems. In the case of TL-Ising model, the behavior of $b_{n}$ defined by the Wightman inner product displays linear growth after a logarithmic correction. The slope $A$ converges to a specific value at infinite temperature, suggesting that the chaotic nature of this spin system is comparatively subdued compared to the GOE.
@@
Note Added. When finishing this manuscript, we notice a new preprint discussing Krylov complexity in RMT at finite temperature has also been proposed Tang (2023). The overlap part is consistent with ours.